# Already Moderate Population Sizes Provably Yield Strong Robustness to Noise

### v3

@@ 5 Runtime Analysis
$\Delta^{+}(d)\geq p_{\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\rfloor}\mathopen{}\mathclose{{}\left\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}}\right\rfloor\geq\frac{1}{4}\mathopen{}\mathclose{{}\left\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}}\right\rfloor.$
@@
$p_{1}\geq\frac{dr}{n}\mathopen{}\mathclose{{}\left(1-\frac{r}{n}}\right)^{n-1}=\frac{dr}{n}\mathopen{}\mathclose{{}\left(\frac{1}{e^{r}}-o(1)}\right)\geq\frac{r}{\lambda e^{r}}(1-o(1))$
@@
$\Delta^{+}(d)$ $\geq 1\cdot\Pr\mathopen{}\mathclose{{}\left[d-\tilde{d}_{y}\geq 1}\right]\geq 1-(1-p_{1})^{\lambda}$
$\geq\frac{1}{2}\min\mathopen{}\mathclose{{}\left(1,\lambda p_{1}}\right)\geq\frac{r}{2e^{r}}(1-o(1))$
@@
$p_{1}\geq\frac{dr}{n}\mathopen{}\mathclose{{}\left(1-\frac{r}{n}}\right)^{n-1}=\frac{dr}{ne^{r}}(1-o(1)).$
@@
$\Delta^{+}(d)$ $\geq\frac{1}{2}\min\mathopen{}\mathclose{{}\left(1,\lambda p_{1}}\right)\geq\frac{\lambda dr}{2ne^{r}}(1-o(1))$
@@
$\sum_{j=1}^{n-d}\Pr[d-\tilde{d}_{y}\geq j]=\sum_{j=1}^{n-d}\mathopen{}\mathclose{{}\left(\Pr[d-\tilde{d}^{\prime}\geq j]}\right)^{\lambda}.$
Let $p_{j}\coloneqq\Pr[d-\tilde{d}^{\prime}\geq j]$. Since $p_{1}$ is at most the probability that at least one 1-bit in $x$ is flipped, we have $p_{1}\leq 1-\mathopen{}\mathclose{{}\left(1-\frac{r}{n}}\right)^{n-d}\leq 1-e^{-r}+o(1)$ (the last inequality is by Lemma 3). Hence,
$p_{1}^{\lambda}\leq\mathopen{}\mathclose{{}\left(1-e^{-r}+o(1)}\right)^{\lambda}\leq\frac{1}{n^{2}}$
@@
$\Delta^{-}(d)=\sum_{j=1}^{n-d}\mathopen{}\mathclose{{}\left(\Pr[d-\tilde{d}^{\prime}\geq j]}\right)^{\lambda}=\sum_{j=1}^{n-d}p_{j}^{\lambda}\leq\frac{n-d}{n^{2}}\leq\frac{1}{n}.$
With Lemma 8 and and with estimates given in Lemmas 9-12 we are now in position to prove our main result, Theorem 7.
@@
Let $h^{+}(d)$ be the lower bound on $\Delta^{+}_{t}(d)$ derived in Lemmas 9-11, that is,
$\Delta^{+}(d)$ $\geq h^{+}(d)\coloneqq\begin{cases}\frac{1}{4}\mathopen{}\mathclose{{}\left\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}}\right\rfloor&\text{ if }d>\frac{n}{\ln n},\\ \frac{r}{2e^{r}}(1-o(1))&\text{ if }d\in\mathopen{}\mathclose{{}\left]\frac{n}{\lambda},\frac{n}{\ln\lambda}}\right],\\ \frac{\lambda dr}{2ne^{r}}(1-o(1))&\text{ if }d\leq\frac{n}{\lambda}.\end{cases}$ (5)
@@
$\Delta^{\text{com}}(d)$ $\geq h^{\text{com}}(d)$
$\coloneqq(1-o(1))\frac{\chi}{q+\chi}\mathopen{}\mathclose{{}\left(h^{+}(d)-h^{-}(d)}\right)-\frac{q\chi}{n}-O\mathopen{}\mathclose{{}\left(\frac{1}{n^{2}}}\right),$
$\Delta^{\text{plus}}(d)$ $\geq h^{\text{plus}}(d)$
$\coloneqq(1-o(1))\frac{\chi}{q+\chi}\mathopen{}\mathclose{{}\left((1-o(1))e^{-q}h^{+}(d)-h^{-}(d)}\right)-\frac{q\chi}{n}$
$-O\mathopen{}\mathclose{{}\left(\frac{1}{n^{2}}}\right).$
We now aim at showing that the first term (the one containing $h^{+}(d)$) is dominant in both $h^{\text{com}}(d)$ and $h^{\text{plus}}(d)$ for all values of $d$. For this we note that by Lemmas 9-11, we have $h^{+}(d)=\omega(\frac{1}{n})$. By Lemma 12 we have $h^{-}(d)=O(\frac{1}{n})=o(h^{+}(d))$. By the constraints $\chi=\Theta(1)$ and $q=O(1)$ we also have $\frac{\chi}{q+\chi}=\Theta(1)$ and $\frac{q\chi}{n}=O(\frac{1}{n})=o(h^{+}(d))$, and $e^{-q}=\Theta(1)$. Therefore, the expressions for the lower bounds on the drift are simplified as follows.
@@
$E[T^{\text{plus}}_{I}]$ $\leq\sum_{d=1}^{n}\frac{1}{h^{\text{plus}}(d)}=\frac{(1+o(1))(q+\chi)e^{q}}{\chi}\sum_{d=1}^{n}\frac{1}{h^{+}(d)},\text{ and}$
$E[T^{\text{com}}_{I}]$ $\leq\sum_{d=1}^{n}\frac{1}{h^{\text{com}}(d)}=\frac{(1+o(1))(q+\chi)}{\chi}\sum_{d=1}^{n}\frac{1}{h^{+}(d)}.$
@@
$\leq n\cdot\frac{9\ln\ln\lambda}{\ln\lambda}+\frac{n}{\ln\lambda}\cdot\frac{2e^{r}}{r}+\frac{2ne^{r}}{\lambda r}\mathopen{}\mathclose{{}\left(\ln\frac{n}{\lambda}+1}\right)$
$=O\mathopen{}\mathclose{{}\left(\frac{n\log\log\lambda}{\log\lambda}}\right)+O\mathopen{}\mathclose{{}\left(\frac{n}{\log\lambda}}\right)+O\mathopen{}\mathclose{{}\left(\frac{n}{\lambda}\log\frac{n}{\lambda}}\right)$
$=O\mathopen{}\mathclose{{}\left(\frac{n}{\lambda}\mathopen{}\mathclose{{}\left(\log\frac{n}{\lambda}+\frac{\lambda\log\log\lambda}{\log\lambda}}\right)}\right).$
@@
$E[T^{\text{plus}}_{I}],E[T^{\text{com}}_{I}]$ $=O\mathopen{}\mathclose{{}\left(\frac{n}{\lambda}\mathopen{}\mathclose{{}\left(\log\frac{n}{\lambda}+\frac{\lambda\log\log\lambda}{\log\lambda}}\right)}\right).$
Since in each iteration of the $(1,\lambda)$ EA we make $\lambda$ fitness evaluations and in each iteration of the $(1+\lambda)$ EA we make $\lambda+1=\Theta(\lambda)$ fitness evaluations, the expected number of fitness evaluations until we find the optimum is at most
$E[T^{\text{plus}}_{F}],E[T^{\text{com}}_{F}]=O\mathopen{}\mathclose{{}\left(n\log\frac{n}{\lambda}+n\frac{\lambda\log\log\lambda}{\log\lambda}}\right).$
We also note that for all $\lambda=O(\frac{\log(n)\log\log(n)}{\log\log\log(n)})$ the first term of the upper bound in Theorem 7 is dominating, hence the expected runtime is $O(n\log(n))$, that is, the same as in the noiseless case.

### v4

@@ 5 Runtime Analysis
$\Delta^{+}(d)\geq p_{\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\rfloor}\left\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\right\rfloor\geq\frac{1}{4}\left\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\right\rfloor.$
@@
$p_{1}\geq\frac{dr}{n}\left(1-\frac{r}{n}\right)^{n-1}=\frac{dr}{n}\left(\frac{1}{e^{r}}-o(1)\right)\geq\frac{r}{\lambda e^{r}}(1-o(1))$
@@
$\Delta^{+}(d)$ $\geq 1\cdot\Pr\left[d-\tilde{d}_{y}\geq 1\right]\geq 1-(1-p_{1})^{\lambda}$
$\geq\frac{1}{2}\min\left(1,\lambda p_{1}\right)\geq\frac{r}{2e^{r}}(1-o(1))$
@@
$p_{1}\geq\frac{dr}{n}\left(1-\frac{r}{n}\right)^{n-1}=\frac{dr}{ne^{r}}(1-o(1)).$
@@
$\Delta^{+}(d)$ $\geq\frac{1}{2}\min\left(1,\lambda p_{1}\right)\geq\frac{\lambda dr}{2ne^{r}}(1-o(1))$
@@
$\sum_{j=1}^{n-d}\Pr[d-\tilde{d}_{y}\geq j]=\sum_{j=1}^{n-d}\left(\Pr[d-\tilde{d}^{\prime}\geq j]\right)^{\lambda}.$
Let $p_{j}\coloneqq\Pr[d-\tilde{d}^{\prime}\geq j]$. Since $p_{1}$ is at most the probability that at least one 1-bit in $x$ is flipped, we have $p_{1}\leq 1-\left(1-\frac{r}{n}\right)^{n-d}\leq 1-e^{-r}+o(1)$ (the last inequality is by Lemma 3). Hence,
$p_{1}^{\lambda}\leq\left(1-e^{-r}+o(1)\right)^{\lambda}\leq\frac{1}{n^{2}}$
@@
$\Delta^{-}(d)=\sum_{j=1}^{n-d}\left(\Pr[d-\tilde{d}^{\prime}\geq j]\right)^{\lambda}=\sum_{j=1}^{n-d}p_{j}^{\lambda}\leq\frac{n-d}{n^{2}}\leq\frac{1}{n}.$
With Lemma 8 and with estimates given in Lemmas 9 to 12 we are now in the position to prove our main result, Theorem 7.
@@
Let $h^{+}(d)$ be the lower bound on $\Delta^{+}_{t}(d)$ derived in Lemmas 9 to 11, that is,
$\Delta^{+}(d)$ $\geq h^{+}(d)\coloneqq\begin{cases}\frac{1}{4}\left\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\right\rfloor&\text{ if }d>\frac{n}{\ln n},\\ \frac{r}{2e^{r}}(1-o(1))&\text{ if }d\in\left]\frac{n}{\lambda},\frac{n}{\ln\lambda}\right],\\ \frac{\lambda dr}{2ne^{r}}(1-o(1))&\text{ if }d\leq\frac{n}{\lambda}.\end{cases}$ (5)
@@
$\Delta^{\text{com}}(d)$ $\geq h^{\text{com}}(d)\coloneqq(1-o(1))\frac{\chi}{q+\chi}\left(h^{+}(d)-h^{-}(d)\right)-\frac{q\chi}{n}-O\left(\frac{1}{n^{2}}\right),$
$\Delta^{\text{plus}}(d)$ $\geq h^{\text{plus}}(d)\coloneqq(1-o(1))\frac{\chi}{q+\chi}\left((1-o(1))e^{-q}h^{+}(d)-h^{-}(d)\right)-\frac{q\chi}{n}-O\left(\frac{1}{n^{2}}\right).$
We now aim at showing that the first term (the one containing $h^{+}(d)$) is dominant in both $h^{\text{com}}(d)$ and $h^{\text{plus}}(d)$ for all values of $d$. For this we note that by Lemmas 9 to 11, we have $h^{+}(d)=\omega(\frac{1}{n})$. By Lemma 12 we have $h^{-}(d)=O(\frac{1}{n})=o(h^{+}(d))$. By the constraints $\chi=\Theta(1)$ and $q=O(1)$ we also have $\frac{\chi}{q+\chi}=\Theta(1)$ and $\frac{q\chi}{n}=O(\frac{1}{n})=o(h^{+}(d))$, and $e^{-q}=\Theta(1)$. Therefore, the expressions for the lower bounds on the drift are simplified as follows.
@@
$E[T^{\text{com}}_{I}]$ $\leq\sum_{d=1}^{n}\frac{1}{h^{\text{com}}(d)}=\frac{(1+o(1))(q+\chi)}{\chi}\sum_{d=1}^{n}\frac{1}{h^{+}(d)},$
$E[T^{\text{plus}}_{I}]$ $\leq\sum_{d=1}^{n}\frac{1}{h^{\text{plus}}(d)}=\frac{(1+o(1))(q+\chi)e^{q}}{\chi}\sum_{d=1}^{n}\frac{1}{h^{+}(d)}.$
@@
$\leq n\cdot\frac{9\ln\ln\lambda}{\ln\lambda}+\frac{n}{\ln\lambda}\cdot\frac{2e^{r}}{r}+\frac{2ne^{r}}{\lambda r}\left(\ln\frac{n}{\lambda}+1\right)$
$=O\left(\frac{n\log\log\lambda}{\log\lambda}\right)+O\left(\frac{n}{\log\lambda}\right)+O\left(\frac{n}{\lambda}\log\frac{n}{\lambda}\right)$
$=O\left(\frac{n}{\lambda}\left(\log\frac{n}{\lambda}+\frac{\lambda\log\log\lambda}{\log\lambda}\right)\right).$
@@
$E[T^{\text{plus}}_{I}],E[T^{\text{com}}_{I}]$ $=O\left(\frac{n}{\lambda}\left(\log\frac{n}{\lambda}+\frac{\lambda\log\log\lambda}{\log\lambda}\right)\right).$
Since each iteration of the $(1,\lambda)$ EA consists of $\lambda$ fitness evaluations, and each iteration of the $(1+\lambda)$ EA consists of $\lambda+1=\Theta(\lambda)$ fitness evaluations, the expected number of fitness evaluations until we find the optimum is at most
$E[T^{\text{plus}}_{F}],E[T^{\text{com}}_{F}]=O\left(n\log\frac{n}{\lambda}+n\frac{\lambda\log\log\lambda}{\log\lambda}\right).$
We note that for all $\lambda=O(\frac{\log(n)\log\log(n)}{\log\log\log(n)})$ the first term of the upper bound in Theorem 7 is dominating, hence the expected runtime is $O(n\log(n))$, that is, the same as in the noiseless case.